OpenStudy (bpaquin07):

Solve with substitution Method: x - 2y = 1 5x-4y=-1

Parth (parthkohli):

start with adding 2y to both sides in the first equation. \( \color{Black}{\Rightarrow x = 2y + 1 }\) Now just substitute it in the second one. \( \color{Black}{\Rightarrow5(2y + 1) -4y = -1 }\) \( \color{Black}{\Rightarrow 10y + 5 - 4y = -1 }\) \( \color{Black}{\Rightarrow 6y + 5 = -1 }\) \( \color{Black}{\Rightarrow 6y = -4 }\) \( \color{Black}{\Rightarrow y = -{4 \over 6} }\) x is two times y plus 1. \( \color{Black}{\Rightarrow x = 2({-2 \over 3}) + 1 }\)